OpenStudy (anonymous):

The cook needs to make the dough for 9 small pizzas. Each pizza dough needs 1 1/3 cup of flour. How much flour does the cook need to make all the pizza dough? Express your answer is simplest form.

OpenStudy (anonymous):

multiply

OpenStudy (boldjon):

got u bro http://www.finecooking.com/recipes/easy_pizza_dough.aspx

OpenStudy (anonymous):

\[9\times 1\tfrac{1}{3}\]

OpenStudy (anonymous):

is it 1 1/27

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what is \(9\times 1\)?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

it is \(9\times 1+9\times \frac{1}{3}\)

OpenStudy (anonymous):

so the answer is 9 1/3

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what is \(9\times \frac{1}{3}\)?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

i don't believe you i am sure you can divide 9 by 3

OpenStudy (anonymous):

its 3

OpenStudy (anonymous):

yes it is and finally, what is \(9+3\)?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

really?

OpenStudy (boldjon):

yes

OpenStudy (anonymous):

im just playing around its 12

OpenStudy (anonymous):

do you know how to multiply a mixed number by a whole number (in general)?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

then you best stay in school

OpenStudy (anonymous):

The answer is 12. The simplest form would be 1 1/3 x 9 = 12 12 bags of flour.. 12 lbs.. Idk.. You should learn